### v1

# ReFeree: Radar-based efficient global descriptor using a Feature and Free space for Place Recognition

Abstract
Radar is highlighted for robust sensing capabilities in adverse weather conditions (e.g. dense fog, heavy rain, or snowfall). In addition, Radar can cover wide areas and penetrate small particles. Despite these advantages, Radar-based place recognition remains in the early stages compared to other sensors due to its unique characteristics such as low resolution, and significant noise. In this paper, we propose a Radar-based place recognition utilizing a descriptor called ReFeree using a feature and free space. Unlike traditional methods, we overwhelmingly summarize the Radar image (i.e. 361.5KB $\rightarrow$ 448B). Despite being lightweight, it contains semi-metric information and is also outstanding from the perspective of place recognition performance. For concrete validation, we test a single session from the MulRan dataset and a multi-session from the Oxford Radar RobotCar and the Boreas dataset. The supplementary materials of our descriptor are available at https://sites.google.com/view/radar-referee.
I Introduction
Place recognition is essential to autonomous vehicle missions (e.g. SLAM (SLAM) or safety navigation). Also, accurate place recognition enables correcting accumulated drift errors in a consistent map. However, as the weather deteriorates, the decline in place recognition performance of most sensors becomes inevitable.
Instead of diverse ranging sensors such as Sonar (Sonar) [1] or LiDAR (LiDAR) [2], we propose a Radar (Radar)-based place recognition called Referee using a feature and free space. Radar can overcome the harsh weather and boast a robust perception performance. In addition, Radar can penetrate small particles, and sensor data is impervious to occlusion. Recently, Radar-based place recognition studies have shown remarkable results [3, 4], but reliable localization remains challenging in various situations. Also, the low resolution and significant noise in sensor data need to be solved.
The main contributions of this work are the following:
  • Effective Description: We propose an efficient three-step description. First, we utilize prominent features from Radar image to minimize noise. Second, we check for areas where features have not been extracted, namely free space (i.e. empty space). Finally, we incorporate free space along the range axis to generate a translational invariant 1D vector. Therefore, the elements forming each vector encapsulate the number of pixels excluding strong signals. Also, since we use the Radar image without transforming it into different properties, the proposed method includes the semi-metric information.
  • Lightweight and Fast: Our descriptor is overwhelmingly lightweight compared to other descriptors. As shown in Fig. 1, our descriptor demonstrates superior performance relative to its size. Moreover, the processing time (generation time + loop searching time) is also short.
  • Various Validation and Open-Source: We conduct experiments in single-session, and multi-session to evaluate diverse and challenging scenarios [5, 6, 7]. Additionally, we aim to contribute to the advancement of the Radar robotics community by sharing our code.
Figure 1: Area Under the ROC Curve (AUC) score per descriptor’s weight in single-session (S) [5] and multi-session datasets (M) [5, 6, 7]. We compare the proposed method with sonar context [1], which utilizes Sonar image similar in form to the Radar image, Raplace [3] and Radar scan context [5] used in Radar.
II Related Works
II-A Place Recognition for Ranging sensors
II-A1 Place Recognition for Sonar
Underwater research is at an earlier stage compared to Radar sensors, but some studies have shown notable results. Lee et al. [8] utilizes Forward Looking Sonar (FLS) to detect artificial markers, thereby enhancing localization performance. However, a drawback of this research is its limitation to place recognition only in man-made environments. Recently, Kim et al. [1] proposes a sonar context that leverages structural information in the underwater environment. This study focuses on understanding the meaningful information from the max intensity of Sonar images and generating global descriptors similar to max pooling. The proposed global descriptor has been found to assist in compensating for accumulated errors in IMU (IMU) data.
II-A2 Place Recognition for LiDAR
LiDAR-based place recognition research is actively progressing in the robotics community. Scan context [2] that utilizes structural information such as building height has shown powerful performance in urban environments. The Scan context of 2D images contains semi-metric information and achieves rotational invariance. The subsequent research algorithm, Scan context++ [9], further enhances performance by achieving translational invariance in addition to rotational invariance. Recently, deep-learning-based research is gaining prominence. Logg3d-NET [10] leverages local features from LiDAR to represent places more prominently. These local features assist in creating better models in the training process. Also, Overlaptransformer [11] utilizes range images in the training process to generate rotational invariant descriptors.
II-B Place Recognition for Radar
Radar-based place recognition is at an incipient stage, but some studies have shown remarkable results [3, 4, 5]. Kim et al. [5] proposed the resized Scan context to transform the radar image itself into a descriptor called Radar scan context. This algorithm also achieves rotational invariance, aiding the ICP (ICP) algorithm with a semi-metric initial pose. Jang et al. [3] utilizes the Radon Transform (RT) to generate rotational and translational invariant descriptors, Raplace. Raplace has contributed to laying the foundation for Radar place recognition. Open-RadVLAD [4] is a vector of locally aggregated descriptors that achieve rotational invariance. This algorithm has demonstrated its performance through an extensive amount of exhaustive results.
III Proposed Method
As described in Fig. 2, our descriptor follows a simple pipeline. The conversion of polar images from radar sensors into cartesian coordinate images is unnecessary. Hence, techniques such as backward warping are not essential, thereby streamlining the description process.
Figure 2: Our proposed pipeline. The navy color represents data, and the blue color represents the algorithm. After acquiring Radar polar images from the sensor, we first proceed with feature extraction and free space counting processes to obtain the final descriptor called Referee. Second, we put the proposed descriptor in the database. Finally, we build a KD-Tree to recognize revisited places through fast searching.
III-A Feature Extraction
Radar image’s intensity called RCS (RCS) is a metric used to indicate how prominently an object appears in radar signal data. RCS varies depending on the target, RCS holds significant importance or structural information. Therefore, how RCS is handled from the Radar’s polar image forms the primary component of the description. Radar image define ×$\mathcal{R}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}}$ in polar coordinates, where $\mathcal{H}$ and $\mathcal{W}$ are the image’s height and width, respectively. The dimensions $\mathcal{H}$ and $\mathcal{W}$ are associated with angle and range, and encapsulating semi-metric information.
We select the feature extraction algorithm proposed in [12] to distinguish valid signals from $\mathcal{R}$. This algorithm involves decomposing the signal into high or low-frequency signals, integrating them according to a Gaussian scaling factor, and thresholding to extract valid features from the raw radar signal. By using this method, we reduce noise from the signal and get higher clarity in classifying areas whether detected objects or vacancy. We define the feature extracted Radar image as ×$\mathcal{R}_{\mathcal{F}}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}}$.
III-B Free space
Although feature extraction reduces noise and leaves only the relatively robust intensity, the features of Radar sensors still have poor resolution, both geometrically and in terms of intensity, compared to denser sensors such as LiDAR. Inspired by [13], this paper proposes a descriptor based on the vacant area (i.e. free-space) in $\mathcal{R}_{\mathcal{F}}$. Free space is efficient due to its conciseness. In addition, it can describe the characteristics of a given area consistently across different sensors. Thus, the descriptor matching based on the overlap between free spaces is robust to the noisy and poor resolution characteristics of the Radar. By improving the methodology of [13] and embedding it as a high-dimensional vector, we reorganize it into a suitable representation for Radar-based place recognition.
Figure 3: Matching graph at the recall position where precision is highest and a Precision-Recall (PR) curve in single session datasets. In the matching graph, the green color represents true loops located within 20m.
III-C ReFeree
Using the features and the free space, we propose a Radar-based efficient global descriptor called ReFeree. To transform $\mathcal{R}_{\mathcal{F}}$ into a ReFeree $\mathbf{K}$, which is an $\alpha$-dimensional vector that each of the elements represents a compression of range bin $\beta$’s free space information. $\mathbf{K}$ is following as:
$\mathbf{K}=\{K_{1},\cdots,K_{\alpha}\}$ (1)
where $\beta\in\mathbb{R}^{\mathcal{H}_{\beta}\times\mathcal{W}}$ is subsection that compose of $\mathcal{R}$, and $\mathcal{H}_{\beta}$ is $\mathcal{H}$ divided by $\alpha$. Free space density of each range bin $K_{\beta}$ is computed as:
$K_{i}={\sum_{j=1}^{\mathcal{\mathcal{H}_{\beta}}}\left(\sum_{k=1}^{r_{j}}s(b_{jk})\right)\over{\mathcal{H}_{\beta}\cdot\mathcal{W}}}$ (2)
where $r$ is element of farthest feature index vector $R=\{r_{1},\cdots,r_{\mathcal{H}_{\beta}}\}$, and it is matched with each row of $\beta$. The notation $b_{jk}$ is a state of single cell located at $j^{th}$ row, $k^{th}$ col in $\beta$, can be converted to numeric value by state function $s(b)$:
$s(b)=\begin{cases}1,&\mbox{if }b\mbox{ is a {free-space}}\\ 0,&\mbox{otherwise}\end{cases}$ (3)
III-D Place Retrieval
For retrieving place from candidates $\mathcal{C}$, we need to establish similarity distance and searching algorithm. Given that ReFeree $\mathbf{K}$ is an $\alpha$-dimensional vector, the KD-Tree algorithm is well-suited for our descriptor, considering the efficiency and speed in searching for distances and optimal candidate $c^{*}$. From this algorithm, we can compute similarity distance $d_{s}$ in the descriptor space, and transitional distance $d_{t}$ using the position of the $c^{*}$. Both distances can be calculated using the L2-norm formula and have similarity threshold $\tau_{s}$ and translational threshold $\tau_{s}$, respectively. Whole retrieval algorithm for query descriptor $\mathbf{K}^{q}$ can be formulated as:
$c^{*}=\underset{c_{k}\in\mathcal{C}}{\operatorname*{argmin}}\;\mathbf{D}(\mathbf{K}^{q},\mathbf{K}^{c_{k}})\;\;s.t.\;\;\mathbf{D}<\mathbf{T}$ (4)
where $\mathbf{D}$ is a set of distance, $\mathbf{D}=\{d_{s},d_{t}\}$, and $\mathbf{T}$ is a set of distance threshold, $\mathbf{T}=\{\tau_{s},\tau_{t}\}$. Only pairs that satisfy both thresholds respectively are treated as detected loops.
IV Experimental Results
We validate our descriptor Referee in various datasets. MulRan [5] includes a variety of sequences. MulRan-DDC provides structural diversity, while MulRan-KAIST can test various dynamic objects and reverse loops. MulRan-Riverside has two bridges, which structurally repeat in parts, and the large-scale MulRan-Sejong is suitable for evaluating multi-session scenarios. Oxford Radar RobotCar [6] and Boreas [7] are datasets acquired over extended periods of time, capturing a wide range of weather conditions. Oxford Radar RobotCar contains some loops, making it suitable for testing single-session scenarios, but it is more appropriate for multi-session evaluations. On the other hand, the Boreas lacks loops, allowing only multi-session evaluations. Additionally, we consider the sonar context applied to Sonar, the radar scan context used for LiDAR, and the Raplace created for Radar as comparative references. All experiments are conducted using Python code on an Intel i7-12700 KF with RTX 3070Ti.
Figure 4: Matching graph at the recall position where precision is highest and a Precision-Recall (PR) curve in multi-session datasets. In the matching graph, the green color represents true loops located within 20m.
IV-A Descriptor Comparison
We compare each descriptor’s size. The proposed method is significantly lighter, ranging from as little as $\times$7 lighter to as much as $\times$70 lighter compared to other methods. Also, the more compressed the current scene, the longer the processing time (i.e. average descriptor generation time + average loop searching time in DDC 01) is, and our descriptor is faster than other methods in terms of time.
TABLE I: Desriptor Comparision
Ours Raplace Scan Context Sonar Context
Descriptor size [B] 448 3008 4928 33728
Processing time [s] 0.106 0.605 0.159 0.125
IV-B Single Session Validation
As depicted in Fig. 3, our method maintains a higher precision for longer than other methods when the recall increases in DDC. As the number of identified true loops increases, we also detect more false loops. AS demonstrated by the PR curve, the proposed method offers superior classification of false loops. Likewise, when recall is 1 at KAIST, the performance across all methods appears comparable. However, considering Riverside where precision is higher, it becomes evident that our approach exhibits exceptional ability in distinguishing false loops.
IV-C Multi Session Validation
Similar to a single session, overall PR-curve performance is all competitive, but when recall is 1, precision is highest as shown in Fig. 4. However, in the 3D matching graph, our performance seems to be poorer than the scan context or sonar context. Because, when the precision is highest, it can be seen that there is little matching in the 3D matching graph because recall is slightly less.
V Conclusion
We propose a lightweight and efficient Radar-based global descriptor called Referee. We validate the superiority of our method across datasets from various scenarios. In future work, we plan to enhance our descriptor for SLAM by incorporating rotation invariance.